# Strong-field Gravitational Wave Lensing in the Kerr Background

### v1

@@ Abstract
Gravitational lensing of gravitational waves (GWs) can encode valuable information about the properties of the intervening lens, yet most existing studies remain restricted to the small-deflection, weak-field regime. To bridge this crucial gap, this work presents the first systematic analysis of strong-field, wave-optical GW lensing by a Kerr black hole (BH), extending recent results for non-rotating lenses obtained in Ref. [Chan:2025wgz] to the astrophysically more relevant case of spinning lens. Using the Mano-Suzuki-Takasugi formalism, we compute the strong-field scattering factor (SFSF) and show that the the spin produces characteristic modifications to the lensed waveform, and high-frequency incident radiation is not strongly absorbed by the BH lens, contrary to earlier claims in Ref. [Chan:2025wgz]. We further derive an explicit expression for the observed waveform for the general source-lens-observer configuration, showcasing the distortions produced by the scattering and quantifying their departure from the Schwarzschild case. Specializing to on-axis scattering, a mismatch analysis for a GW150914-like source lensed by a Kerr BH of mass $M=10^{2}M_{\odot}$ situated $100GM/c^{2}$ away from the source reveals percentage-level deviations from the direct (unscattered) wave at scattering angles near $30^{\circ}$, across a range of lens spin values. The mismatch exhibits a sharp decrease at larger scattering angles. For a fixed scattering angle, however, the mismatch shows only a weak dependence on the BH spin in the case of on-axis scattering, which may improve for more general off-axis considerations. The framework developed here offers a unified treatment of strong-field GW scattering in Kerr spacetime and provides tools for interpreting future high-precision observations of compact-object lenses in the wave-optics regime.
@@ I Introduction
Since the first direct detection of gravitational waves (GWs) with GW150914 [LIGOScientific:2016aoc], identifying GW signals from compact binary coalescences (CBCs) with the Laser Interferometer Gravitational-wave Observatory (LIGO) [LIGOScientific:2014pky] and Virgo [VIRGO:2014yos] has become a routine. With more than $200$ confident detections to date [LIGOScientific:2018mvr, LIGOScientific:2020ibl, LIGOScientific:2021usb, KAGRA:2021vkt, LIGOScientific:2025slb], we are now able to probe the most extreme and dynamic regions of spacetime like never before [LIGOScientific:2016lio, LIGOScientific:2017zic, LIGOScientific:2019fpa, LIGOScientific:2021sio, LIGOScientific:2025obp, Siegel:2025xgb]. The ever-growing catalog of CBCs has also enabled population studies of stellar-mass black holes (BHs) [LIGOScientific:2020kqk, KAGRA:2021duu, LIGOScientific:2025pvj], revealing several surprising features, such as lower/upper mass-gap events [Mehta:2021fgz, Zevin:2020gma, Edelman_2021], high-mass and misaligned-spin BHs [Pierra:2024fbl, LIGOScientific:2025pvj], the presence of hybrid binaries (e.g., BH-neutron star systems) [Chattopadhyay:2022cnp, Qin:2024ojw], and indications of hierarchical mergers [Li:2022gly, Tagawa:2021ofj]. The GW detection rate, currently several events per month, is expected to increase significantly as the sensitivities of second-generation ground-based detectors, including KAGRA [Somiya:2011np, Aso:2013eba, KAGRA:2018plz, KAGRA:2020tym], improve and the addition of LIGO-India [LigoIndia, Unnikrishnan:2013qwa, Ajith:2024inj] strengthens the global network. The growth will be further accelerated with the advent of next-generation observatories such as the Einstein Telescope (ET) [Punturo:2010zz, Hild:2010id], Cosmic Explorer (CE) [Reitze:2019iox, LIGOScientific:2016wof, Regimbau:2016ike], and the space-based Laser Interferometer Space Antenna (LISA) [LISA:2017]. With these advancements, a host of subtle phenomena are expected to become observable. Among these, GW lensing [1936Sci....84..506E, PhysRev.51.290, 1992grle.book.....S] stands out as a particularly promising probe, allowing us to explore structures across a wide range of astrophysical and cosmological scales.
@@
Gravitational lensing of GWs arise when spacetime curvature produced by an intervening mass alters the propagation of the wavefront [1971NCimB...6..225L, Ohanian:1974ys]. In the limit where the GW wavelength $\lambda$ is much shorter than the characteristic lens scale (i.e., $\lambda\ll 2GM/c^{2}$, where $M$ is the lens mass), gravitational lensing can be described using ray optics analogous to the standard electromagnetic case, resulting in multiple images of the same source that arrive at the detector with characteristic time delays and relative magnifications [Bernardeau:1999mh, Takahashi:2016jom, Dai:2017huk, Smith:2017jdz, Liao:2017ioi, Haris:2018vmn, Li:2018prc, Broadhurst:2018saj, Broadhurst:2019ijv, Broadhurst:2020cvm, Dai:2020tpj, Ezquiaga:2020gdt, Caliskan:2022wbh, Vijaykumar:2022dlp, Barsode:2024zwv]. However, when the GW wavelength becomes comparable to or larger than the Schwarzschild radius of the lens (i.e., $\lambda\gtrsim 2GM/c^{2}$), the geometric-optics approximation breaks down, necessitating a full wave-optics treatment [1986ApJ...307...30D, Nakamura:1997sw, Nakamura:1999uwi, 2003ApJ...595.1039T]. In this regime, interference and diffraction effects become significant, producing characteristic frequency-dependent modulations in the observed GW amplitude and phase, which can bias the estimated parameters or other analyses such as tests of general relativity [Jung:2017flg, Shan:2020esq, Mishra:2021xzz, Meena:2022unp, Mishra:2023ddt, Mishra:2023vzo, Cremonese:2021puh, Cremonese:2021ahz, Shan:2023ngi, Rao:2025poe]. For current ground-based detectors, an intervening compact object with mass $\sim 10-10^{5}$~M can induce observable wave-optics signatures. For space-based detectors such as LISA, which are sensitive to much lower GW frequencies (and hence probe much larger lens masses), wave-optics effects are expected to be particularly prominent [Caliskan:2022hbu, Pijnenburg:2024btj]. GW lensing therefore naturally bridges two complementary regimes: the geometric-optics limit relevant for galaxy- and cluster-scale lenses, and the wave-optics limit essential for compact-object lenses such as stellar-mass and supermassive black holes (SMBHs).
@@
Driven by these exciting prospects, the theoretical study of GW lensing has advanced considerably in recent years. Early studies primarily focused on point-mass and galactic-scale lenses, employing diffraction integrals and Kirchhoff-type formalisms to compute magnifications and time delays within the wave-optics regime [Ohanian:1974ys, 1986ApJ...307...30D, 1992grle.book.....S, Nakamura:1997sw, Baraldo:1999ny, 2003ApJ...595.1039T, Biesiada:2021pzo, Dalang:2021qhu, Tambalo:2022plm, Harikumar:2023gzh, Leung:2023lmq, Deka:2024ecp]. These approaches successfully captured the interference patterns characteristic of microlensed GWs and were later generalized to include realistic lens mass profiles, dark matter substructures, and cosmological corrections [Mishra:2021xzz, Meena:2022unp, Villarrubia-Rojo:2024xcj, Deka:2025vzx, Shan:2022xfx, Shan:2023ngi, Liu:2023ikc]. More recently, investigations of strong-field lensing by Schwarzschild BHs [Chan:2025wgz] in the long-wavelength regime have revealed a wealth of wave-optical phenomena, such as frequency-dependent amplification, intricate phase structures, and polarization effects absent in the geometric optic limit. However, this study is restricted to non-rotating lenses, leaving open the important question of how BH spin, one of the defining features of astrophysical BHs, affects GW propagation in the strong-field regime.
@@
In particular, for the special case of on-axis scattering, a mismatch analysis for a GW150914-like event shows a percentage-level deviations from the direct wave at scattering angle near $30^{\circ}$, which falls sharply at larger scattering angles. Here, we must emphasize that a percentage-level mismatch lies close to the detectability threshold $\sim(\text{signal-to-noise ratio})^{-2}$ (which is around $0.1-1\%$ for typical LIGO SNRs) of the current ground-based detectors. By contrast, next-generation detectors will drive this threshold down to $10^{-4}-10^{-6}$ level, making a percentage-level or even smaller mismatches unambiguously detectable. Interestingly, at a fixed scattering angle, the mismatch for on-axis scattering shows only a mild sensitivity to the lens spin parameter, which may improve for more general off-axis scattering. Although our formulation is valid for off-axis scattering as well, it posses significant numerical challenges, which we aim to address in a dedicated future work. These results demonstrate that strong-field wave-optical lensing by Kerr BHs can generate potentially detectable distortions in GW signals and should therefore be accounted for in future searches targeting compact-object lensing in dense astrophysical environments.
@@
Astrophysically, such strong-field lensing effects may arise in several realistic scenarios. For example, CBCs formed via dynamical formation channels in dense globular clusters [Rodriguez:2018rmd, Antonini:2020xnd, Ubach:2025dob] or active galactic nuclei (AGN) disks [2012MNRAS.425..460M, Tagawa:2019osr, Mckernan:2017ssq, Grobner:2020drr, Ubach:2025dob] may experience strong-field Kerr lensing of their emitted GWs by a central SMBH, where the spin of the lens plays a decisive role. Another promising setting involves hierarchical triple systems [Antonini:2013tea, Silsbee:2016djf, Fragione:2019zhm, Liu:2020gif, Ubach:2025dob], in which an inner compact binary inspirals within the gravitational potential of an outer spinning BH. In such configurations, the emitted GWs can undergo strong-field Kerr lensing as they propagate through the curved spacetime near the outer companion. Although the overall occurrence rates of these scenarios remain uncertain, even a small fraction of detectable mergers (on the order of a few percent) occurring in such environments could provide valuable probes of gravity in the strong and dynamical regimes predicted by GR. Indeed, events such as GW190521, which has been suggested to originate within an AGN disk [Graham:2020gwr, Morton:2023wxg, Leong:2024nnx], demonstrate that such environments may already be within observational reach of current ground-based detectors. With the advent of next-generation observatories offering enhanced sensitivity at lower frequencies, the prospects for detecting strong-field lensing signatures in GWs will only grow more promising.
@@ II GWs Scattering by Kerr Lens
In order to investigate the gravitational lensing of GWs by the strong field of a Kerr BH, let us first setup a few basic notations. The source of GWs is considered to be a coalescing compact binary, while the lens is taken to be a rotating Kerr BH characterized by its mass $M$ and spin angular momentum $J=a\,M=\chi\,M^{2}$. We assume the source is sufficiently far away from the lens such that, at the lens location, the incident GWs can be well described as plane waves. We also have $r_{\rm SL}\gg R_{L}$, where $r_{\rm SL}$ is the physical lens-source distance and $R_{L}=2M$ is the Schwarzschild radius of the lens. For reasons that will be clear later, we shall primarily focus on configurations in which the source and the lens are relatively nearby, i.e., $r_{\rm SL}\ll r_{\rm SO}\sim r_{\rm LO}$, with $r_{\rm LO}$ ($r_{\rm SL}$) being the physical observer-lens (lens-source) distance. Under this assumption, the apparent angular position of the observer with respect to the lens, $\Omega_{\rm LO}=\{\theta_{\rm LO},\phi_{\rm LO}\}$, can be taken to be approximately the same as that with respect to the source, $\Omega_{\rm SO}=\{\theta_{\rm SO},\phi_{\rm SO}\}$. While treating the scattering problem, we shall use the lens-centred coordinates $x^{\mu}=\{t,x,y,z\}$ (unit vectors along the spatial directions will be denoted with hats) and their spherical polar analogues, with the $z$-axis chosen as the lens’ spin axis. The observer’s polar coordinates are then denoted as $\vec{r}_{\rm LO}\equiv(r_{\rm LO},\theta_{\rm LO}=\theta_{o},\phi_{\rm LO}=\phi_{o})$. The setup is schematically illustrated in Fig. 1.
Figure 1: The source (S) is a coalescing binary, which is situated close to a lensing BH (L) of mass $M$ and spin-angular momentum $J$. A portion of the radiation emitted by the source is scattered ($h^{\mu\nu}_{\rm SW}$) by the lens in the direction of the observer (O). The coordinate system is aligned so that the spin of the lensing BH is along the $z$-axis. The x-axis is chosen such that the source and lens lie on the xzx-z plane. The angular coordinates of the observer/detector (O) are denoted as $(\theta_{o},\phi_{o})$. The scattered wave $h^{\mu\nu}_{\rm SW}$, at the observer can be computed in the framework of BH perturbation theory (BHPT). The radiation from the source can also reach the observer “directly” as $h^{\mu\nu}_{\rm dir}$. The direct component is only weakly affected by the lens for large $\theta_{o}\sim 30^{\circ}$.

### v2

# Strong-field Gravitational-Wave Lensing in the Kerr Background

@@ Abstract
Gravitational lensing of gravitational waves (GWs) can encode valuable information about the properties of the intervening lens, but most existing studies remain restricted to the small-deflection, weak-field regime. To bridge this crucial gap, this work presents the first systematic analysis of strong-field, wave-optical GW lensing by a Kerr black hole (BH), extending recent results obtained in Ref. [Chan:2025wgz] for non-rotating lens to the astrophysically more relevant case of spinning-lens. Using the Mano-Suzuki-Takasugi formalism, we define and compute the strong-field scattering factor (SFSF) and show that the the spin produces characteristic modifications to the lensed waveform, and high-frequency incident radiation is not strongly absorbed by the BH lens, contrary to earlier claims in Ref. [Chan:2025wgz]. We further derive an explicit expression for the observed waveform for the general source-lens-observer configuration, showcasing the distortions produced by the scattering and quantifying their departure from the Schwarzschild case. Specializing to on-axis scattering, a mismatch analysis for a GW150914-like source lensed by a Kerr BH of mass $M=10^{2}~\mathrm{M}_{\odot}$ situated $100GM/c^{2}$ away from the source reveals percentage-level deviations from the direct (unscattered) wave at scattering angles near $30^{\circ}$, across a range of lens spin values. The mismatch generally decreases as the scattering angle increases, but this behavior can change substantially when polarization mixing induced by scattering becomes significant. In such cases, components that are absent or suppressed in the direct signal may become appreciable once scattering effects are taken into account. For a fixed scattering angle, however, the mismatch shows only a weak dependence on the BH spin in the case of on-axis scattering, which may improve for more general off-axis considerations. The framework developed here offers a unified treatment of strong-field GW scattering in Kerr spacetime and provides tools for interpreting future high-precision observations of compact-object lenses in the wave-optics regime.
@@ I Introduction
Since the first direct detection of gravitational waves (GWs) with GW150914 [LIGOScientific:2016aoc], identifying GW signals from compact binary coalescences (CBCs) with the LIGO [LIGOScientific:2014pky] and Virgo [VIRGO:2014yos] has become a routine. With more than $200$ confident detections to date [LIGOScientific:2018mvr, LIGOScientific:2020ibl, LIGOScientific:2021usb, KAGRA:2021vkt, LIGOScientific:2025slb], we are now able to probe the most extreme and dynamic regions of spacetime like never before [LIGOScientific:2016lio, LIGOScientific:2017zic, LIGOScientific:2019fpa, LIGOScientific:2021sio, LIGOScientific:2025obp, Siegel:2025xgb]. The ever-growing catalog of CBCs has also enabled population studies of stellar-mass black holes (BHs) [LIGOScientific:2020kqk, KAGRA:2021duu, LIGOScientific:2025pvj], revealing several surprising features, such as lower/upper mass-gap events [Mehta:2021fgz, Zevin:2020gma, Edelman_2021], high-mass and misaligned-spin BHs [Pierra:2024fbl, LIGOScientific:2025pvj], the presence of hybrid binaries (e.g., BH-neutron star systems) [Chattopadhyay:2022cnp, Qin:2024ojw], and indications of hierarchical mergers [Li:2022gly, Tagawa:2021ofj]. The GW detection rate, currently several events per month, is expected to increase significantly as the sensitivities of second-generation ground-based detectors, including KAGRA [Somiya:2011np, Aso:2013eba, KAGRA:2018plz, KAGRA:2020tym], improve and the addition of LIGO-India [LigoIndia, Unnikrishnan:2013qwa, Ajith:2024inj] strengthens the global network. The growth will be further accelerated with the advent of next-generation observatories such as the Einstein Telescope (ET) [Punturo:2010zz, Hild:2010id], Cosmic Explorer (CE) [Reitze:2019iox, LIGOScientific:2016wof, Regimbau:2016ike], and the space-based Laser Interferometer Space Antenna (LISA) [LISA:2017]. With these advancements, a host of subtle phenomena are expected to become observable. Among these, GW lensing [1936Sci....84..506E, PhysRev.51.290, 1992grle.book.....S] stands out as a particularly promising probe, allowing us to explore structures across a wide range of astrophysical and cosmological scales.
@@
Gravitational lensing of GWs arise when spacetime curvature produced by an intervening mass alters the propagation of the wavefront [1971NCimB...6..225L, Ohanian:1974ys]. In the limit where the GW wavelength $\lambda$ is much shorter than the characteristic lens scale (i.e., $\lambda\ll 2GM/c^{2}$, where $M$ is the lens mass), gravitational lensing can be described using ray optics analogous to the standard electromagnetic case, resulting in multiple images of the same source that arrive at the detector with characteristic time delays and relative magnifications [Bernardeau:1999mh, Takahashi:2016jom, Dai:2017huk, Smith:2017jdz, Liao:2017ioi, Haris:2018vmn, Li:2018prc, Broadhurst:2018saj, Broadhurst:2019ijv, Broadhurst:2020cvm, Dai:2020tpj, Ezquiaga:2020gdt, Caliskan:2022wbh, Vijaykumar:2022dlp, Barsode:2024zwv]. However, when the GW wavelength becomes comparable to or larger than the Schwarzschild radius of the lens (i.e., $\lambda\gtrsim 2GM/c^{2}$), the geometric-optics approximation breaks down, necessitating a full wave-optics treatment [1986ApJ...307...30D, Nakamura:1997sw, Nakamura:1999uwi, 2003ApJ...595.1039T]. In this regime, interference and diffraction effects become significant, producing characteristic frequency-dependent modulations in the observed GW amplitude and phase, which can bias the estimated parameters or other analyses such as tests of general relativity [Jung:2017flg, Shan:2020esq, Mishra:2021xzz, Cremonese:2021puh, Cremonese:2021ahz, Meena:2022unp, Shan:2023ngi, Mishra:2023ddt, Mishra:2023vzo, Liu:2023emk, Liu:2024xxn, Lo:2024wqm, Ezquiaga:2025gkd, Rao:2025poe]. For current ground-based detectors, an intervening compact object with mass $\sim 10-10^{5}$~M can induce observable wave-optics signatures. For space-based detectors such as LISA, which are sensitive to much lower GW frequencies (and hence probe much larger lens masses), wave-optics effects are expected to be particularly prominent [Caliskan:2022hbu, Pijnenburg:2024btj]. GW lensing therefore naturally bridges two complementary regimes: the geometric-optics limit relevant for galaxy- and cluster-scale lenses, and the wave-optics limit essential for compact-object lenses such as stellar-mass and supermassive black holes (SMBHs).
@@
Driven by these exciting prospects, the theoretical study of GW lensing has advanced considerably in recent years. Early studies primarily focused on point-mass and galactic-scale lenses, employing diffraction integrals and Kirchhoff-type formalisms to compute magnifications and time delays within the wave-optics regime [Ohanian:1974ys, 1986ApJ...307...30D, 1992grle.book.....S, Nakamura:1997sw, Baraldo:1999ny, 2003ApJ...595.1039T, Biesiada:2021pzo, Dalang:2021qhu, Tambalo:2022plm, Harikumar:2023gzh, Leung:2023lmq, Deka:2024ecp]. These approaches successfully captured the interference patterns characteristic of microlensed GWs and were later generalized to include realistic lens mass profiles, dark matter substructures, and cosmological corrections [Mishra:2021xzz, Meena:2022unp, Shan:2022xfx, Shan:2023ngi, Liu:2023ikc, Villarrubia-Rojo:2024xcj, Deka:2025vzx, Vujeva:2025nwg, Vujeva:2025kko, Bulashenko:2025vdx]. More recently, investigations of strong-field lensing by Schwarzschild BHs [Chan:2025wgz] in the long-wavelength regime have revealed a wealth of wave-optical phenomena, such as frequency-dependent amplification, intricate phase structures, and polarization effects absent in the geometric optic limit. However, this study is restricted to non-rotating lenses, leaving open the important question of how BH spin, one of the defining features of astrophysical BHs, affects GW propagation in the strong-field regime.
@@
In particular, for the special case of on-axis scattering, our mismatch analysis for a GW150914-like event (i.e., a binary with intrinsic parameters similar to those of GW150914) shows a percentage-level deviations from the direct wave at scattering angle near $30^{\circ}$, which falls sharply at larger scattering angles. Here, we must emphasize that a percentage-level mismatch lies close to the detectability threshold $\sim(\text{SNR})^{-2}$, which is around $0.1-1\%$ for typical (and conservative) SNRs in the range of 103010-30 for the current ground-based detectors [Lindblom:2008cm, Damour:2010zb, Cornish:2011ys, Thompson:2025hhc]. By contrast, next-generation detectors will drive this threshold down to $10^{-4}-10^{-6}$ level, making a percentage-level or even smaller mismatches unambiguously detectable [Thompson:2025hhc]. Interestingly, at a fixed scattering angle, the mismatch for on-axis scattering shows only a mild sensitivity to the lens spin parameter, which may improve for more general off-axis scattering. Although our formulation is valid for off-axis scattering as well, it posses significant numerical challenges, which we aim to address in a dedicated future work. These results demonstrate that strong-field wave-optical lensing by Kerr BHs can generate potentially detectable distortions in GW signals and should therefore be accounted for in future searches targeting compact-object lensing in dense astrophysical environments.
@@
Astrophysically, such strong-field lensing effects may arise in several realistic scenarios. For example, CBCs formed via dynamical formation channels in dense globular clusters [Rodriguez:2018rmd, Antonini:2020xnd, Ubach:2025dob] or active galactic nuclei (AGN) disks [2012MNRAS.425..460M, Tagawa:2019osr, Mckernan:2017ssq, Grobner:2020drr, Ubach:2025dob] may experience strong-field Kerr lensing of their emitted GWs by a central SMBH, where the spin of the lens plays a decisive role. A recent analysis of gravitational lensing of GWs emitted by a stellar-mass compact binary orbiting a spinning SMBH is presented in Ref. [Santos:2025ass]. Another promising setting involves hierarchical triple systems [Antonini:2013tea, Silsbee:2016djf, Fragione:2019zhm, Liu:2020gif, Ubach:2025dob], in which an inner compact binary inspirals within the gravitational potential of an outer spinning BH. In such configurations, the emitted GWs can undergo strong-field Kerr lensing as they propagate through the curved spacetime near the outer companion. Although the overall occurrence rates of these scenarios remain uncertain, even a small fraction of detectable mergers (on the order of a few percent) occurring in such environments could provide valuable probes of gravity in the strong and dynamical regimes predicted by GR. Indeed, events such as GW190521 (the most massive binary reported in GWTC-3), which has been suggested to originate within an AGN disk [Graham:2020gwr, Morton:2023wxg, Leong:2024nnx], demonstrate that such environments may already be within observational reach of current ground-based detectors. With the advent of next-generation observatories offering enhanced sensitivity at lower frequencies, the prospects for detecting strong-field lensing signatures in GWs will only grow more promising.
@@ II GWs Scattering by Kerr Lens
In order to investigate the gravitational lensing of GWs by the strong field of a Kerr BH, let us first setup a few basic notations. The source of GWs is considered to be a coalescing compact binary, while the lens is taken to be a rotating Kerr BH characterized by its mass $M$ and spin angular momentum $J=a\,M=\chi\,M^{2}$. We assume the source is sufficiently far away from the lens such that, at the lens location, the incident GWs can be well described as plane waves. We also have $r_{\rm SL}\gg R_{L}$, where $r_{\rm SL}$ is the physical lens-source distance and $R_{L}=2M$ is the Schwarzschild radius of the lens. For reasons that will be clear later, we shall primarily focus on configurations in which the source and the lens are relatively nearby, i.e., $r_{\rm SL}\ll r_{\rm SO}\sim r_{\rm LO}$, with $r_{\rm LO}$ ($r_{\rm SO}$) being the physical observer-lens (observer-source) distance. Under this assumption, the apparent angular position of the observer with respect to the lens, $\Omega_{\rm LO}=\{\theta_{\rm LO},\phi_{\rm LO}\}$, can be taken to be approximately the same as that with respect to the source, $\Omega_{\rm SO}=\{\theta_{\rm SO},\phi_{\rm SO}\}$. While treating the scattering problem, we shall use the lens-centred coordinates $x^{\mu}=\{t,x,y,z\}$ (unit vectors along the spatial directions will be denoted with hats) and their spherical polar analogues, with the $z$-axis chosen as the lens’ spin axis. The observer’s polar coordinates are then denoted as $\vec{r}_{\rm LO}\equiv(r_{\rm LO},\theta_{\rm LO}=\theta_{o},\phi_{\rm LO}=\phi_{o})$. The setup is schematically illustrated in Fig. 1.
Figure 1: The source (S) is a coalescing binary, which is situated close to a lensing BH (L) of mass $M$ and spin-angular momentum $J$. A portion of the radiation emitted by the source is scattered ($h^{\mu\nu}_{\rm SW}$) by the lens in the direction of the observer (O). The coordinate system is aligned so that the spin of the lensing BH is along the $z$-axis. The $x$-axis is chosen such that the source and lens lie on the xzxz-plane. The angular coordinates of the observer/detector (O) are denoted as $(\theta_{o},\phi_{o})$. The scattered wave $h^{\mu\nu}_{\rm SW}$, at the observer can be computed in the framework of BH perturbation theory (BHPT). The radiation from the source can also reach the observer “directly” as $h^{\mu\nu}_{\rm dir}$. The direct component is only weakly affected by the lens for large $\theta_{o}\sim 30^{\circ}$.